OpenStudy (anonymous):

Which of the following equations has no solution? #1: 3x + 1 = 2x + 1 #2: 3x + 1 = 3x + 1 #3: 3x + 1 = 3x + 2

OpenStudy (shamil98):

option 3

OpenStudy (shamil98):

3x + 1 = 3x + 2 -3x -3x \[1\]

OpenStudy (shamil98):

\[1\neq2\]

OpenStudy (anonymous):

Whenever you solve an equation which ends up in an untrue statement, the equation has no solution. For example, x + 1 = x + 2 If you solve that equation by subtracting x from both sides, you get 1 = 2 which is an untrue statement. Therefore the equation x + 1 = x + 2 has no solutions. Now use this concept and try to answer your question.

OpenStudy (anonymous):

thanks...both of you